# DAWN: Domain-Adaptive Weakly Supervised Nuclei Segmentation via Cross-Task Interactions

### v1

@@ III-C Combined Pseudo-label Optimization
Due to domain shift resulting from nuclei morphology and size variations across different organs, there may be an under-segmentation problem in the target domain prediction $\hat{y}_{t}^{seg}$. To solve the problem, we devise a combined pseudo-label (CPL) optimization method by combining the outputs of the segmentation and detection networks. This method utilizes the superior object localization capabilities of the detection network and effectively compensates for the flaw of the segmentation network. The schematic diagram of the method is depicted in Fig. 4.
The CPL method comprises two key components: mask fusion and distance filter. In mask fusion, the prediction outputs from both networks are combined to achieve comprehensive nuclei segmentation. The distance filter is then applied to refine the fused results by eliminating incorrectly segmented nuclei. The resulting outputs from the CPL serve as the ground truth annotations for supervising the subsequent training.
@@ III-C1 Mask Fusion
Based on the prediction output $\hat{q}_{t}^{det}$ of the detection network, we employ a threshold filter to preserve the high confidence region of prediction $\hat{q}_{t}^{det}$ as shown in Fig. 4 (a) and obtain the initialized detection $M_{det}$ as shown in Fig. 4 (b), which represents the nuclei region:
$M_{det}=I_{[\hat{q}_{t}^{det}>\theta]}\in\{0,1\}^{N\times N},$ (7)
where $\hat{q}_{t}^{det}$ represents the probability output of detection model, $\theta$ is the threshold value, $M_{det}$ represents generated nuclei region. At the same time, we transform the probability output $\hat{p}_{t}^{seg}$ of the segmentation network, obtaining the initialized $M_{ini}$, which has the under-segmentation problem as shown in Fig. 4 (c). Then, we calculate a union set for $M_{ini}$ and $M_{det}$ as shown in Fig. 4 (d):
@@ III-D Interactive Supervision Training Method
This sub-section proposes a dynamic pseudo-labels optimization method to facilitate domain adaptation in the segmentation network. Specifically, the pseudo-labels generated from the CPL module in the $(i-1)^{th}$ round will be viewed as supervision information for the training of segmentation and detection networks in the $i^{th}$ round. By continuously optimizing the segmentation network, this method helps the progressive domain transfer of the network. Our interactive supervision process is shown in Fig. 3 (c)-(d) and Algorithm 1 Line 13-16.
@@
$L_{dyn}=CE(M_{pse},\hat{p}_{t}^{seg})+MSE(M_{pse},\hat{q}_{t}^{det}),$ (9)
where $M_{pse}$ is the pseudo-label generated by the CPL module, $\hat{p}_{t}^{seg}$ and $\hat{q}_{t}^{det}$ represent the predicted output of segmentation and detection networks in the target domain.
@@ IV-A Datasets
In this subsection, we briefly introduce the PanNuke, MoNuSeg used in the source domain (♣), and TNBC, CryoNuSeg, ConSeP used in the target domain (♠) in this paper. The PanNuke, MoNuSeg, TNBC, CryoNuSeg, and ConSeP datasets are sampled at ×40 magnification. The lizard dataset is sampled at ×20 magnification.
@@ IV-B Implementation Details and Evaluation Metrics
Our DAWN uses a U-Net framework with a ResNet [64] backbone as a detection network. Meantime, a ResNet encoder combined with a DenseNet decoder is used as a segmentation network. Besides, all experiments are run with PyTorch on an Nvidia RTX 3090 GPU. We set loss weights $\alpha=0.1$, $\beta=0.15$ used in Eq. 1. The hyperparameters used Eq. 4 and Eq. 7 are shown in Table I. $r_{1}$ is set according to the mean nuclei radius of the validation set, and $r_{2}$ is set to double $r_{1}$ for balancing foreground and background pixels. $\theta$ is set according to the detection network output on the validation set. We evaluate our experiments over metrics of DICE, AJI[60], DQ[65], SQ[65] and PQ[65]. Among those, AJI, DQ, SQ, and PQ are instance-level metrics. DQ evaluates the detection quality, SQ evaluates segmentation quality, and PQ is the product of the DQ and SQ. Throughout all the tables in this paper, we bold the best and underline the second best.
@@
TNBC PseEdgeNet[66] 0.682 0.497 0.688 0.547 0.394 Lizard PseEdgeNet [66] 0.545 0.140 0.229 0.534 0.105
(a) WeakSAM[67] 0.834 0.697 0.795 0.782 0.620 (c) WeakSAM[67] 0.643 0.279 0.314 0.608 0.267
@@
CryoNuSeg PseEdgeNet[66] 0.620 0.321 0.507 0.529 0.306 ConSeP PseEdgeNet [66] 0.569 0.226 0.245 0.307 0.171
(b) WeakSAM[67] 0.754 0.480 0.627 0.685 0.449 (d) WeakSAM[67] 0.669 0.378 0.384 0.680 0.401
@@ IV-C Comparison with Weakly Supervised Methods
We compare our proposed domain-adaptive weakly supervised nuclei segmentation method with seven other models on four datasets, and the comparison methods consist of DDTNet[14], BoNuS[15], WeakSeg[16], MaskGANet[45], C2FNet[13], PseEdgeNet[66], and WeakSAM[67]. The experiment results are shown in Table II. In the table, the best weakly supervised method is bold with blue text, and the second best weakly supervised method is underlined with blue text. For fair comparison, for the two-stage methods WeakSeg, BoNuS and DDTNet, we use the same pseudo labels as our DAWN. In detail, the pre-trained segmentation network is test on target dataset and the prediction outputs is used as pseudo labels. In the experiment, we use PanNuke as source domain dataset.
@@ IV-C1 Comparisons on Breast Cancer TNBC
We first validate our proposed domain-adaptive weakly supervised nuclei segmentation on a breast cancer dataset, and the experimental results are shown in Table II. From this table, we observe that our proposed method outperforms existing state-of-the-art weakly supervised segmentation methods. Compared with the best method, our DAWN achieves improvements of 2.2% and 1.6% in terms of DICE and AJI metrics, respectively. Moreover, our method exhibits significant advantages over traditional two-stage weakly supervised models such as MaskGANet and WeakSeg. Our method surpasses MaskGANet by at least 10% and WeakSeg by at least 4.0% on all five metrics.
Furthermore, we conduct visualization comparison experiments as shown in the first two rows of Fig. 5. From the figure, it can be observed that our proposed DAWN, WeakSeg, DDTNet, and BoNuS can effectively segment nuclei. However, methods like MaskGANet and PENet exhibit weaker morphological learning capabilities. We attribute this to pre-trained segmentation networks, where the original domain’s segmentation network contains a wealth of prior morphological information that improves nuclei segmentation performance.
@@ IV-C2 Comparisons on Pan-Cancer CryoNuSeg
We evaluate the performance of our proposed domain adaptation method for weakly supervised segmentation on a challenging pan-cancer dataset called CryoNuSeg. It is worth noting that the CryoNuSeg dataset exhibits significant variations in nuclei morphology, making it a problematic dataset for training. Our experimental results are presented in Table II (b). From these results, it can be observed that our proposed DAWN outperforms the best-performing method, WeakSAM, by 5.0% and 2.8% in terms of DICE and AJI metrics, respectively. Furthermore, compared to the second-best method, DDTNet, our DAWN achieves approximately 10% improvement in DICE, AJI, and PQ metrics.
We perform visual comparison experiments, and the results are shown in the third and fourth rows of Fig. 5. The results of WeakSeg, MaskGANet, C2FNet, and PENet show that these methods exhibit poorer segmentation results when faced with significant morphological variations of nuclei. In contrast, our method still achieves outstanding performance. We attribute this phenomenon to the fact that two-stage methods struggle to adapt their pseudo-labels to the variations in staining and morphology, leading to lower segmentation performance. However, our method utilizes a domain transfer strategy that leverages prior knowledge from the source domain and introduces a progressive domain adaptation method to adapt to the target domain’s image characteristics, resulting in better segmentation results.
@@ IV-C3 Comparisons on Colorectal Cancer Lizard and ConSeP
We further validate the performance of our proposed method on colorectal cancer pathological images. From the Lizard results in Table II (c), we observe that existing weakly supervised methods perform poorly, especially C2FNet, MaskGANet, and PseEdgeNet, with PQ scores of only 0.134, 0.163, and 0.105, respectively. The phenomenon is because the Lizard dataset is sampled at the magnification of 20, resulting in densely distributed and small-sized nuclei, posing challenges for the segmentation task. Nevertheless, our method still achieves improvements compared to the best-performing WeakSeg. We observe improvements of 1.4% and 2.2% for DICE and AJI metrics, respectively. Additionally, from the fifth and sixth rows of Fig. 5, it can be seen that the nuclei in the Lizard dataset are small and exhibit dense distribution, and most models struggle to produce satisfactory results. However, our approach can better identify the nuclei compared to other methods.
From the ConSeP results in Table II (d), it can be observed that our method achieves significant improvements compared to other weakly supervised models. Compared to C2FNet, our DAWN model achieves improvements of 5.6% and 4.5% in terms of DICE and AJI, respectively. From the visual results in the last two rows of Fig. 5, it can be seen that the nuclei in the ConSeP dataset often have blurry morphology and dense distribution. Other methods struggle to identify the boundaries and distinguish individual nuclei instances accurately. In contrast, our method partially addresses this issue and somewhat improves the segmentation performance.
@@ IV-D Comparison with Fully Supervised Methods
To compare the performance difference between our proposed weakly supervised segmentation method and existing fully supervised segmentation methods, we compare our DAWN with six state-of-the-art models, including U-Net [8], HoverNet [30], CDNet [32], DCAN [33], Micro-Net [68], and Dist [61]. In the experiment, we use PanNuke as the source domain and TNBC and ConSeP as the target domains. For a fair comparison, all comparison models are pre-trained in PanNuke and then fine-tuned in the four datasets. The results are shown in Table II. In the table, the best fully supervised method is bold with red text, and the second best fully supervised method is underlined with red text.
@@ IV-D1 Quantitative Comparisons
From Table II, it can be seen that our method achieves comparable performance to fully supervised methods on the TNBC dataset. In the DICE, our method outperforms the best-performing CDNet by 0.4% and significantly surpasses other traditional supervised methods such as U-Net and DCAN. On the CryoNuSeg and ConSeP datasets, although our method does not achieve the best performance, it is only 1.8% and 0.4% lower than the second-best HoverNet method in terms of AJI. However, our method still performs poorly on the Lizard dataset. While it outperforms U-Net and DCAN, there is still a performance gap compared to other methods. We attribute this phenomenon to the limitations of our domain adaptation method when dealing with densely distributed nuclei data, and using point annotations for domain transfer still poses challenges.
@@ IV-D2 Qualitative Comparisons
Furthermore, we provide comparison visualization in Fig. 6. From the figure we can see our weakly supervised segmentation method achieves comparable performance with supervised methods. Surprisingly, our method even outperforms Micro-Net and U-Net. This indicates an accurate pseudo label initialization and an elaborate pseudo label optimization also provide fine supervision information although without pixel-level mask.
@@
IV-E Domain Adaptation Evaluation
We conduct a domain adaptation comparison with three domain adaptation methods: SIFA[17], PDAM[18] and SFDA-CellSeg[69]. The comparison experiment is designed from three perspectives.
IV-E1 PanNuke as Source dataset
To compare the different domain adaptation methods, we use a large pan-cancer dataset, PanNuke, as the source domain data and TNBC (breast cancer) and CryoNuSeg (pan-cancer) as the target domain data for the experiment. The experiment results are shown in Table III. From this table, we can observe that our method achieves significant improvements both on TNBC and CryoNuSeg datasets. Regarding the AJI metric, our method outperforms the second-best method by 4.63% and 4.49% on TNBC and CryoNuSeg, respectively.
TABLE III: The domain adaptation comparison experiments from source domain PanNuke to target domain TNBC and CryoNuSeg.
@@ IV-E1 PanNuke as Source dataset
TABLE IV: The domain adaptation comparison experiments from source domain MoNuSeg to target domain TNBC and CryoNuSeg.
@@ IV-E2 MoNuSeg as Source dataset
To validate the influences of data scale, we conducted domain adaptation experiments on a smaller pan-cancer dataset, MoNuSeg. The experiment results are shown in Table IV. When using MoNuSeg as the source domain data, the AJI metric of our method surpasses the second-best method by 3.36% and 3.24% on TNBC and CryoNuSeg target data, respectively. This indicates that our method is effective on small-scale datasets. Besides, comparing the results in Table III and Table IV, the latter’s performance significantly decreases. On the CryoNuSeg data, the DICE metric for the three comparison methods decreased by 0.33%, 2.06%, and 2.25%, while our proposed method decreased by only 1.77%, showing a smaller decline compared to the PDAM and SFDA-CellSeg. The phenomenon suggests that our method is less affected by the data scale disturbance.
IV-E3 Cross-cancer Domain Adaptation
We conduct cross-cancer domain adaptation experiments on the ConSeP (colon cancer) and TNBC (breast cancer) datasets, and the experiment results are shown in Table V. We can observe that our method achieves the best performance among these methods. Particularly on the TNBC, the AJI surpasses the second-best method by 4.20%. These experiments demonstrate the significant advantage of our method. Based on this phenomenon, we can conclude that our method is least affected by the source domain data. Our method can still exploit essential prior information for different cancer types to enhance the segmentation performance.
Figure 7: The comparison of before and after domain adaptation. Blue circles indicate instances which are not recognition in source only method; Red and white circles represent accurate instance segmentation in our DAWN; Yellow arrow represents more smoother boundary of our DAWN, which is enlarged in bottom-right corner.
TABLE V: The cross-cancer domain adaptation experiments from source domain ConSeP (TNBC) to target domain TNBC (ConSeP).

### v3

@@ III-C Combined Pseudo-label Optimization
Due to domain shifts caused by variations in nuclei morphology and size across different organs, under-segmentation can occur in the target domain predictions. To address this issue, we propose a Combined Pseudo-Label (CPL) optimization method, which integrates the outputs of both the segmentation and detection networks. This approach utilizes the superior object location capabilities of the detection network to effectively compensate for the shortcomings of the segmentation network. A schematic diagram of the process is provided in Fig. 4.
The CPL method comprises two key components: mask fusion and the distance filter. In mask fusion, the prediction outputs from both networks are combined to achieve comprehensive nuclei segmentation. The distance filter is then applied to refine the fused results by eliminating incorrectly segmented nuclei. The resulting outputs from the CPL serve as ground-truth annotations for supervising the subsequent training.
@@ III-C1 Mask Fusion
Based on the prediction output $\hat{q}_{t}$ of the detection network, we employ a threshold filter to preserve the high confidence region of prediction $\hat{q}_{t}$ as shown in Fig. 4 (a) and obtain the initialized detection $M_{det}$ as shown in Fig. 4 (b), which represents the nuclei region:
$M_{det}=I_{[\hat{q}_{t}>\theta]}\in\{0,1\}^{N\times N},$ (7)
where $\hat{q}_{t}$ represents the probability output of detection model, $\theta$ is the threshold value, $M_{det}$ represents generated nuclei region. At the same time, we transform the probability output $\hat{p}_{t}$ of the segmentation network, obtaining the initialized $M_{ini}$, which has the under-segmentation problem, as shown in Fig. 4 (c). Then, we calculate a union set for $M_{ini}$ and $M_{det}$ as shown in Fig. 4 (d):
@@ III-D Interactive Supervision Training Method
This subsection proposes a dynamic pseudo-label optimization method to facilitate domain adaptation in the segmentation network. Specifically, the pseudo-labels generated from the CPL module in the $(t-1)^{th}$ round serve as supervision information for the training of segmentation and detection networks in the $t^{th}$ round. By iterative optimizing the segmentation network, this method enables progressive domain transfer. The interactive supervision process is shown in Fig. 3 (c)-(d) and Algorithm 1 Line 13-16.
@@
$L_{dyn}=\text{CE}(M_{pse},\hat{p}_{t})+\text{MSE}(M_{pse},\hat{q}_{t}),$ (9)
where $M_{pse}$ is the pseudo-label generated by the CPL module, $\hat{p}_{t}$ and $\hat{q}_{t}$ represent the predicted output of segmentation and detection networks in the target domain.
@@ IV-A Datasets
In this subsection, we introduce the PanNuke, MoNuSeg datasets used in the source domain (♣), and TNBC, CryoNuSeg, ConSeP datasets used in the target domain (♠). In detail, the PanNuke, MoNuSeg, TNBC, CryoNuSeg, and ConSeP datasets are sampled at ×40 magnification. The Lizard dataset is sampled at ×20 magnification.
@@ IV-B Implementation Details and Evaluation Metrics
Our DAWN uses a U-Net framework with a ResNet [71] backbone as a detection network. Meantime, a ResNet encoder combined with a DenseNet decoder is used as a segmentation network. All experiments are run with PyTorch on an Nvidia RTX 3090 GPU.
We set the loss weights in Eq. 1 to $\alpha=0.1$ and $\beta=0.15$. The hyperparameters specified in Eq. 4 and Eq. 7 are detailed in Table I. Specifically, $r_{1}$ and $\sigma$ are determined based on the mean and standard deviation of the nuclear radius in the validation set, with $r_{2}$ set at twice the value of $r_{1}$ to balance the foreground and background pixels. The parameter $\theta$ is established according to the output of the detection network in the validation set, while $d$ is determined empirically. We evaluate our experiments using the metrics of DICE, AJI[67], DQ[72], SQ[72] and PQ[72].
@@
TNBC SPF-TN [44] 0.806 0.687 0.812 0.739 0.611 Lizard SPF-TN[44] 0.610 0.221 0.314 0.602 0.261
(a) PENet [73] 0.682 0.497 0.688 0.547 0.394 (c) PENet [73] 0.545 0.140 0.229 0.534 0.105
WeakSAM[74] 0.834 0.697 0.795 0.782 0.620 WeakSAM[74] 0.643 0.279 0.314 0.608 0.267
@@
CryoNuSeg SPF-TN [44] 0.676 0.449 0.611 0.634 0.391 ConSeP SPF-TN[44] 0.632 0.392 0.373 0.667 0.382
(b) PENet[73] 0.620 0.321 0.507 0.529 0.306 (d) PENet [73] 0.569 0.226 0.245 0.307 0.171
WeakSAM[74] 0.754 0.480 0.627 0.685 0.449 WeakSAM[74] 0.669 0.378 0.384 0.680 0.401
@@ IV-C Comparison with Weakly Supervised Methods
We compare our proposed domain-adaptive weakly supervised nuclei segmentation method with eight other models on four datasets. The comparison methods consist of DDTNet[16], BoNuS[17], WeakSeg[18], MaskGANet[38], C2FNet[15], PENet [73], SPF-TN [44], and WeakSAM[74]. The experiment results are shown in Table II. In the table, the best weakly supervised method is bold, and the second best weakly supervised method is underlined. For a fair comparison, we use the same pseudo labels as our DAWN for the two-stage methods WeakSeg, BoNuS, and DDTNet. In detail, the pre-trained segmentation network is tested on the target dataset, and the prediction outputs are used as pseudo labels. In the experiment, we use PanNuke as the source domain dataset.
@@ IV-C1 Comparisons on Breast Cancer TNBC
We validate our proposed DAWN on a breast cancer dataset, and the experimental results are shown in Table II (a). From this table, we observe that our proposed method outperforms existing state-of-the-art weakly supervised segmentation methods. Compared with the best method, our DAWN achieves improvements of 2.2% and 1.6% in terms of DICE and AJI metrics, respectively. Moreover, our method exhibits significant advantages over traditional two-stage weakly supervised models such as MaskGANet and WeakSeg. In detail, our method surpasses MaskGANet by at least 10% and surpasses WeakSeg by at least 4.0% on all five metrics.
Furthermore, we conduct visualization comparison experiments as shown in the first two rows of Fig. 5. From the segmentation results, it can be seen that our method has clear advantages. On the TNBC and CryoNuSeg datasets, our method segments cell nuclei contours more accurately, while on the Lizard and ConSeP datasets, when the nuclei distribution is dense, our method can more comprehensively identify nuclei.
@@ IV-C2 Comparisons on Pan-Cancer CryoNuSeg
We evaluate the performance of our proposed method, DAWN, on the challenging pan-cancer CryoNuSeg dataset. This dataset presents significant variations in nuclear morphology, posing substantial challenges for training. Our experimental results, as shown in Table II (b), demonstrate that DAWN outperforms the leading method, WeakSAM, with improvements of 5.0% and 2.8% in DICE and AJI metrics, respectively. Additionally, DAWN shows notable gains over the second-best approach, achieving improvements of approximately 10%, 6%, and 8% in DICE, AJI, and PQ metrics, respectively.
We perform visual comparison experiments, and the results are shown in the third and fourth rows of Fig. 5. The segmentation results of WeakSeg, MaskGANet, and C2FNet show that these methods exhibit poorer segmentation results when faced with significant morphological variations of the nuclei. In contrast, our method still achieves outstanding performance. We attribute this phenomenon to the fact that two-stage methods struggle to adapt their pseudo-labels to the variations in staining and morphology, leading to lower segmentation performance. However, our method utilizes a domain transfer strategy that leverages prior knowledge from the source domain and introduces a progressive domain adaptation method to adapt to the target domain’s image characteristics.
@@ IV-C3 Comparisons on Colorectal Cancer Lizard and ConSeP
We further validate the performance of our proposed method in pathological images of colorectal cancer. As shown by the results of the Lizard dataset in Table II (c), existing weakly supervised methods, particularly C2FNet, MaskGANet, and PENet, perform poorly, with PQ scores of 0.134, 0.163, and 0.105, respectively. This underperformance can be attributed to the fact that the Lizard dataset is sampled at a magnification of ×$\times 20$, resulting in densely distributed and small nuclei, which presents significant challenges for segmentation. Despite these difficulties, our method achieves improvements over the best-performing method, WeakSeg, with increases of 1.4% and 2.2% in DICE and AJI metrics, respectively. Furthermore, as illustrated in the fifth and sixth rows of Fig. 5, our method maintains excellent segmentation performance even with smaller nuclei. In contrast, the segmentation results of DDTNet, which performed well on the TNBC and CryoNuSeg datasets, are unsatisfactory on Lizard. This highlights that our method is less affected by the sampling magnification compared to other models.
As shown in the ConSeP results in Table II (d), our proposed DAWN model achieves significant improvements over other weakly supervised models. Specifically, DAWN outperforms C2FNet with gains of 5.6% in DICE and 4.5% in AJI. The visual results in the last two rows of Fig. 5 highlight that the nuclei in the ConSeP dataset often exhibit blurry morphology and are densely distributed, making it difficult for other methods to accurately delineate boundaries and distinguish individual nuclei instances. In contrast, our method addresses these challenges more effectively, leading to improved segmentation performance.
@@ IV-D Comparison with Fully Supervised Methods
To compare the performance difference between our proposed weakly supervised segmentation method and existing fully supervised segmentation methods, we compare our DAWN with seven state-of-the-art models, including U-Net [10], HoverNet [33], CDNet [35], DCAN [32], Micro-Net [75], Dist [68], and AOST[76]. In the experiment, we use PanNuke as the source domain and TNBC and ConSeP as the target domains. For a fair comparison, all comparison models are pre-trained in PanNuke and then fine-tuned in the four datasets. The results are shown in Table II. In the table, the best fully supervised method is bold, and the second best fully supervised method is underlined.
@@ IV-D1 Quantitative Comparisons
As shown in Table II, our method achieves performance comparable to fully supervised methods on the TNBC dataset. Specifically, our method surpasses the best-performing CDNet by 0.9% in DICE and significantly outperforms other traditional supervised methods such as U-Net and DCAN. On the CryoNuSeg and ConSeP datasets, although our method does not achieve the top performance, it is only 1.8% and 0.4% lower than the second-best HoverNet in terms of AJI. However, our method underperforms on the Lizard dataset. While it outperforms U-Net and DCAN, a performance gap remains compared to other methods. We attribute this to the inherent limitations of weakly supervised tasks, as using point annotations for domain transfer continues to present challenges.
@@ IV-D2 Qualitative Comparisons
Furthermore, we provide comparison visualization in Fig. 6. In the figure, blue stands for correctly true positive pixels, red stands for false positive pixels and green stands for false negative pixels. From the figure we can see our weakly supervised segmentation method achieves comparable performance with supervised methods. Surprisingly, our method even outperforms Micro-Net and U-Net. This indicates an accurate pseudo label initialization and an elaborate pseudo label optimization also provide fine supervision information although without pixel-level mask.
@@
TABLE III: The domain adaptation comparison experiments from PanNuke dataset to TNBC and CryoNuSeg datasets.
@@
IV-E Domain Adaptation Evaluation
We conduct a domain adaptation comparison with six domain adaptation methods: SIFA[19], PDAM[20] and SFDA-CellSeg[77], ADPL[53], CyCADA[54], and DDMRL[55]. The domain adaptation experiments are designed from the following three perspectives.
IV-E1 PanNuke as Source dataset
To compare the different domain adaptation methods, we use a large pan-cancer dataset, PanNuke, as the source domain data and TNBC (breast cancer) and CryoNuSeg (pan-cancer) as the target domain data for the experiment. The experiment results are shown in Table III. From this table, we can observe that our method achieves significant improvements both on TNBC and CryoNuSeg datasets. Regarding the AJI metric, our method outperforms the second-best method by 1.58% and 2.78% on TNBC and CryoNuSeg, respectively.
TABLE IV: The domain adaptation comparison experiments from MoNuSeg dataset to TNBC and CryoNuSeg datasets.
@@ IV-E2 MoNuSeg as Source dataset
To validate the influences of data scale, we conducted domain adaptation experiments on a smaller pan-cancer dataset, MoNuSeg. The experiment results are shown in Table IV. When using MoNuSeg as the source domain data, the AJI metric of our method surpasses the second-best method by 2.10% and 1.91% on TNBC and CryoNuSeg target data, respectively. This indicates that our method is effective on small-scale datasets. Besides, comparing the results in Table III and Table IV, the latter’s performance significantly decreases. On the CryoNuSeg data, the DICE of the ADPL and CyCADA decreased by 3.20% and 3.71%, while our proposed method decreased by only 1.77%, showing a smaller decline compared to the ADPL and CyCADA. The phenomenon suggests that our method is less affected by the data scale disturbance.
TABLE V: The cross-cancer domain adaptation experiments from ConSeP (TNBC) dataset to TNBC (ConSeP) dataset.
@@
IV-E3 Cross-cancer Domain Adaptation
We conduct cross-cancer domain adaptation experiments on the ConSeP (colon cancer) and TNBC (breast cancer) datasets, and the experiment results are shown in Table V. We can observe that our method achieves the best performance among these methods. Particularly on the TNBC, the AJI surpasses the second-best method by 4.03%. These experiments demonstrate the significant advantage of our method. Based on this phenomenon, we can conclude that our method is least affected by the source domain data. Our method can still exploit essential prior information for different cancer types to enhance the segmentation performance.